40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 565     Entry time: Wed Jun 25 11:36:14 2008
Author: Max Jones 
Type: Update 
Category: Computer Scripts / Programs 
Subject: First Week Update 
For the first week I have been modifying the noise budget script in caltech/NB to run with 40 m parameters and data. As per Rana's instructions, I have tried to run the script with only seismic and Darm sources. This involves identifying and changing channel names and altering paramter files (such as NB/ReferenceData/C1IFOparams.m). To supply the parameter files, I have copied the H1 files with (as yet only) slight modification. The channel name changes have been made to mirror the sites for the most part. Two figures are attached which show the current noise budget. The Day plot was taken 6/23/08 at ~10:30 am. The night plot was taken 6/22/08 at ~11:00 pm . Note that the SRD curve is for the sites and not for the 40 m (I hope to change that soon). Also in one of the plots the DARM noise signal is visible. Obviously this needs work. A list of current concerns is

1) I am using a seismic transfer function made by previous SURF student Ryan Kinney to operate with channels of the form C1:PEM_ACC-ETMY_Y (should I be using C1:DMF-IX_ACCY?) and the channels I am currently using are the acceleraometers for the mode cleaner with names of the form C1:PEM_ACC-MC1_X. Rana said that he thinks these may be the same but I need to be sure.

2) We don't have a DARM_CTRL channel but the code requires it, currently I am using DARM_ERR as a substitute which is probably partly responsible for the obvious error in DARM noise.

Any suggestions are appreciated. Max.
Attachment 1: C1_NoiseBudgetPlot_Day.eps  25 kB  | Hide | Hide all | Show all
C1_NoiseBudgetPlot_Day.eps
Attachment 2: C1_NoiseBudgetPlot_Night.eps  86 kB  | Show | Hide all | Show all
ELOG V3.1.3-